OpenStudy (anonymous):

find dy/dx of y=lnx-1/e^x+1

OpenStudy (anonymous):

is it: \[y=\ln(x) - 1/(e ^{x}+1)\] ?

myininaya (myininaya):

\[y=\frac{lnx-1}{e^x+1} ?\]

OpenStudy (anonymous):

yep

myininaya (myininaya):

well the way he has it is what you have inink

OpenStudy (anonymous):

yep to which one?

myininaya (myininaya):

say first or second one

OpenStudy (anonymous):

second. thanks guys

myininaya (myininaya):

\[y'=\frac{(lnx-1)'(e^x+1)-(e^x+1)'(lnx-1)}{(e^x+1)^2}\]

myininaya (myininaya):

(lnx-1)'=1/x-0=1/x (e^x+1)=e^x+0=e^x

myininaya (myininaya):

\[y'=\frac{\frac{1}{x}*(e^x+1)-e^x(lnx-1)}{(e^x+1)^2}\]

OpenStudy (anonymous):

\[y'=(1/x *(e ^{x}+1) -e ^{x}(lnx-1))/(e ^{x}+1)^{2}\] see myininaya's ... simplify if needed

OpenStudy (anonymous):

oooo...see i was confused about what to do with the [e ^{x}. thought i had to do somethin tricksy lol. thanks a lot =)

myininaya (myininaya):

\[y'=\frac{x}{x}*\frac{\frac{1}{x}*(e^x+1)-e^x(lnx-1)}{(e^x+1)^2}=\frac{(e^x+1)-xe^x(lnx-1)}{x(e^x+1)^2}\]